You've just joined the investment banking firm of Dewey, Cheatum, and Howe. They've offered you two different salary arrangements. You can have $123,000 per year for the next two years, or you can have $45,000 per year for the next two years, along with a $30,000 signing bonus today. The bonus is paid immediately, and the salary is paid at the end of each year.


Required:

(a)

If the interest rate is 9 percent compounded monthly, what is the present value of the first arrangement?

  (Click to select)   217,411.14   216,370.68   215,258.56   206,648.21   223,868.9
(b)

If the interest rate is 9 percent compounded monthly, what is the present value of the second arrangement?

  (Click to select)   108,753.13   113,103.26   104,403   109,160   109,840.66
Answer #1

-112/ITUDY2) 111 ДA B C D F G EAR 9.38% au WNP Year Cash flow 1 $123,000.00 $123,000.00 $0.00 $112,451.29 $102,807.26 PV 7 8

F16 Α B C D E F G EAR =(1+9%/12)^12-1 O 60 vu WNA a) Year Cash flow PV 1 123000 =E5/(1+$D$2)^E4 2 123000 =F5/(1+$D$2)^F4 =D5/